OpenStudy (anonymous):

A simple pendulum having charge +Q, length l, and mass m is hung from a fixed support between the horizontal plates of a charged capacitor. What is the time period of vibration ?

OpenStudy (anonymous):



OpenStudy (anonymous):

The formula governing the motion of a regular pendulum in a gravitational field is \[ \tau = I \alpha \] \[ -mgL \sin(\theta) = mL^2 \ddot{\theta} \] using the small angle approximation, \[mL^2 \ddot{\theta} + mgL\theta = 0\] or \[\ddot{\theta} + \frac{g}{L} \theta = 0\] which is the equation of motion for a pendulum with angular frequency \[ \omega = \sqrt{g/L} \] and time period \[T = 2\pi \sqrt{\frac{L}{g}} \] In this case, everything is the same except for the torque. The torque is no longer mgL sin(theta).